Length of a rectangle floor is 1m. less than twice its witdth. If a diagonal of the rectangle is 17 meters, find the length and width of the floor. Thank you for your help.
Answer by mananth(16946)

Length of a rectangle floor is 1m. less than twice its witdth. If a diagonal of the rectangle is 17 meters, find the length and width of the floor. Thank you for your help.
...
let width be x
length = 2x-1
...
the sides and the diagonal form a rectangle
x^2+(2x-1)^2=17^2
x^2+4x^2-4x+1=289
5x^2-4x+1-289=0
5x^2-4x-288=0
5x^2-40x+x+36x-288=0
5x(x-8)+36(x-8)=0
(x-8)(5x+36)=0
x= 8m the width
length = 2x-1
=16-1
=15 m the length
...
m.ananth@hotmail.ca